Determine the two ordered pair solutions of the equation x-y=-3 , given (0, ), and (-5, ).
Answer by Fombitz(32388)

Plug x into the equation and solve for y.
I'll do one, you do the other the same way.
x-y=-3
0-y=-3
-y=-3
y=3
(0,3)